In 1931, President Herbert Hoover was paid a salary of $75,000. Government statistics show a consumer price index of 15.2 for 1931 and 214.5 for 2009. President Hoovers 1931 salary was equivalent to a 2009 salary of about: Select one a. $ 1,058,388 b. $1,140,000 c. $15,525,000 d. $ 5,507
Answer #1

Answer

The correct answer is (a) $1,058,388

In 1931, Consumer Price Index = 15.2 and salary = 75000

Hence Real Salary = Salary/ CPI

=> 1931 real Salary(i.e. Salary with price adjusted) = 75000/15.2 = 4934.21

In 2009, Consumer Price Index = 214.5.

We want real Salary in 2009 to be equivalent to real salary in 1931

2009 real Salary(i.e. Salary with price adjusted) = Salary in 2009/214.5

=>  Salary in 2009/214.5 = 75000/15.2

=> Salary in 2009 = $1058388

Hence, the correct answer is (a) $1058388